s.t every pair of lines can be represented by ax^2+2hxy+by^2+k(x^2+y^2) k is const

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akash

ax^2+2hxy+by^2+k(x^2+y^2)=0

(a+k)x2 +2hxy +(b+k)y2 =0

a'x2 +2hxy +b'y2 =0   this is the general form of pair of st line through origin so every line can be represented in above manar


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand -
post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.
 All the best.
 Regards,
Askiitians Experts
Badiuddin